Determine the magnitudes of FA and FB so that the resultant force passes through point O of the column. Suppose that a = 150 mm , b = 180 mm , c = 550 mm , and d = 800 mm .Determine the magnitudes of FA and F so that the resultant force passes through point O of the column. Suppose that a 150 mm
Answer #1

taking moment about x axis, we have

Fa * 0.6 + 8 * d - 6 * c  - Fb * 0.75 = 0

put value of d and c, we have

Fa * 0.6 + 8 * 0.8 - 6 * 0.55 - Fb * 0.75 = 0

0.6Fa - 0.75Fb + 3.1 = 0 ------------ (1)

now, taking moment about y axis

8 * a + Fb * b - 6 * a - Fa * b = 0

put value of a and b , we get

8 * 0.15 + Fb * 0.18 - 6 * 0.15 - Fa * 0.18 = 0

0.18Fb - 0.18 Fa + 0.3 = 0 --------- (2)

solving (1) and (2) , we get

Fa = 29.0 kN

Fb = 27.3 kN